Explain in steps how you would sketch the graph of the following function in the interval from 0 to 2(pi) y = cos 4x
Answer by lwsshak3(11628)

Explain in steps how you would sketch the graph of the following function in the interval from 0 to 2(pi) y = cos 4x
..
Standard form for equation for cosine function: y=Acos(Bx-C), With A=amplitude, period=2π/B, and phase-shift=C/B
..
Given cosine function has an amplitude=1 and period 2π/4=π/2, phase-shift=0
1/4 period=π/8
To graph, make tick marks on the x-axis at 0, π/8, 2π/8, 3π/8 and 4π/8.
You now will have the following points to plot the curve:
(0,1),(π/8,0),(2π/8,-1),(3π/8,0), (4π/8,1) (This covers 1-period, π/2)
Repeat plotting curve 3 more periods on tick marks spaced=1/4 period=π/8 until 2π is reached.